### v1

# Kullback–Leibler Potential for Non-Ergodic Replication Dynamics: An Information-Theoretic Second Law

Abstract
This study aims to quantify and visualize the degradation of fidelity (information degradation) that inevitably accompanies the replication of information within the framework of information thermodynamics and to propose an information-theoretic formulation of the second law based on this phenomenon. While previous research in information thermodynamics has focused on the thermodynamic costs associated with information “erasure” or “measurement” through concepts such as Landauer’s principle and mutual information, little systematic discussion has addressed the inherently irreversible nature of “replication” itself and the accompanying degradation of information structure. In this study, we construct a mathematical model of information replication using a discrete Markov model and Gaussian convolution, and quantify changes in information at each replication step: Shannon entropy, cross-entropy, and the Kullback–Leibler divergence (KLD). The monotonic decrease of KLD exhibits a Lyapunov-like property, which can be interpreted as a potential analogous to the free energy in the process by which a nonequilibrium system converges to a particular steady state. Furthermore, we extend this framework to the potential applicability to biological information processes such as DNA replication, showing that the free energy required for degradation and repair can be expressed in terms of KLD. This contributes to building a unified information-thermodynamic framework for operations such as replication, transmission, and repair of information.
I Introduction
The second law of thermodynamics links irreversibility to entropy production. In information thermodynamics, the physical cost of manipulating information has been clarified for erasure and measurement through Landauer’s principle and its extensions: erasing one bit dissipates at least $k_{\mathrm{B}}T\ln 2$ of heat [1], and feedback/measurement operations admit refined second-law-like inequalities expressed through mutual information [2, 3, 4]. In contrast, the inherently irreversible nature of replication - biological or digital - has not been formulated in a unified information–thermodynamic framework. From a biochemical point of view, kinetic proofreading already indicates that high-fidelity copying requires free energy expenditure [5], yet a coarse-grained information-theoretic law for replication itself has remained elusive.
This work. We model replication as a discrete-time Markov map on probability distributions under block invariance (non-ergodic constraints). Given an initial distribution $p_{0}$, let $\Pi(p_{0})$ denote the reachable steady set composed of blockwise invariant distributions that preserve the initial block masses (formal definitions in Sect. II). Our central object is Kullback–Leibler (KL) potential.
$V(p)\ :=\ \inf_{\pi\in\Pi(p_{0})}D_{\mathrm{KL}}\!\bigl{(}p\,\|\,\pi\bigr{)},$ (1)
which measures the excess information retained beyond the conserved coarse variables. We prove that $V$ is a Lyapunov function for block-invariant replication dynamics $p_{n+1}=Tp_{n}$:
$V(p_{n+1})\ \leq\ V(p_{n})\qquad(n=0,1,2,\dots),$ (2)
with equality only in blockwise steady states (Theorem 1). Moreover, when each block is ergodic with a unique invariant $\pi_{j}^{\ast}$, the potential vanishes asymptotically:
$V(p_{n})\ \xrightarrow[n\to\infty]{}\ 0.$ (3)
Unlike the classical ergodic case—where the KL divergence to a single invariant measure is Lyapunov—our $V$ in (1) is a minimum over steady mixtures constrained by the initial block masses, capturing the essential non-ergodicity of replication.
Instantiations and readouts. We instantiate the framework in two settings: (i) image copying via Gaussian convolution (an information-losing map), and (ii) a block diagonal substitution model for DNA replication with proofreading. In both cases, we report the stepwise evolution of Shannon entropy $H$, cross-entropy $H_{\times}$, KL divergence $D_{\mathrm{KL}}$, and the potential $V$ using the single set of definitions and Theorem 1 from Sect. II. All entropies/divergences are in nats unless noted. Connections to prior information-thermodynamic perspectives on replication and non-equilibrium free energy analogies (e.g., interpreting cross-entropy differences as generated heat) are discussed in the context of our results [19, 20].
II General framework and main theorem
II.1 Block invariance and reachable steady set
The information replication operations we study (image blurring, base substitution with proofreading, etc.) are inherently local: one step only rearranges probability mass within a constrained neighborhood (spatial or compositional), but does not freely mix all microscopic states. As a result, certain coarse observables, e.g., the total mass inside a spatial region of an image, or the adenine(A) -thymine (T) rich block and guanine(G) -cytosine (C) rich block in DNA to be replicated as a template, are effectively conserved by every replication step. When such conserved coarse variables exist, the Markov dynamics does not converge to a single global invariant distribution; instead, the state space decomposes into components that evolve independently. We call this situation non-ergodic in the sense that long-term behavior retains memory of the initial coarse structure through the conserved weights.
(i) In the image model, a blockwise Gaussian convolution forbids smoothing across block boundaries; the total intensity in each block is preserved at every step. (ii) In the DNA model, a block diagonal substitution kernel reflects biochemical or sequence constraints (e.g., AT- vs. GC-rich segments); block masses (AT vs. GC) remain fixed while within-block compositions relax. (iii) More generally, locality, topological separation, or routing restrictions in copying channels induce invariant ’blocks’ that prevent global mixing. Because replication preserves these coarse variables, the appropriate equilibrium notion is not a single point but a set of steady states obtained by fully relaxing within each invariant component while keeping the initial coarse weights fixed. This motivates modeling the state space as a disjoint union of blocks and defining a reachable steady set against which we will measure the distance using a KL-based potential. We consider a finite state space partitioned into disjoint blocks.
$\mathcal{X}\ =\ \bigsqcup_{j=1}^{m}\,\mathcal{X}_{j},$ (4)
and a Markov kernel $T$ acting on $\mathcal{P}(\mathcal{X})$. Here $\mathcal{P}(\mathcal{X})$ denotes the probability simplex over $\mathcal{X}$, i.e. the set of all probability distributions $p:\mathcal{X}\to[0,1]$ with $\sum_{x\in\mathcal{X}}p(x)=1$; we identify $p$ with a row vector and let $T$ act by $(pT)(y)=\sum_{x\in\mathcal{X}}p(x)\,T(x,y)$, where for each $x$ the map $A\mapsto T(x,A)$ is a probability measure on subsets of $\mathcal{X}$ (so, in the finite case, $T$ is a row stochastic matrix). The dynamics is block-invariant if the transitions never cross the block boundaries, i.e.,
$x\in\mathcal{X}_{j}\ \Longrightarrow\ T(x,\mathcal{X}_{j})=1\qquad(j=1,\dots,m).$ (5)
In above, for a Markov kernel $T$, $T(x,A)$ is the one-step probability to land in $A$ given the current state $x$. Thus, for $x\in\mathcal{X}_{j}$, the condition $T(x,\mathcal{X}_{j})=1$ means the next state lies in the same block with probability 1; equivalently $\sum_{y\in\mathcal{X}_{j}}T(x,y)=1$ and $T(x,y)=0$ for all $y\notin\mathcal{X}_{j}$. In matrix form, each row indexed by $x\in\mathcal{X}_{j}$ has the support contained in $\mathcal{X}_{j}$, so $T$ is block-diagonal and block masses are preserved.Block invariance guarantees conservation of block masses $w_{j}$ but not convergence within blocks. Our potential $V(p)=\inf_{\pi\in\Pi(p_{0})}D_{\mathrm{KL}}(p\|\pi)$ over the reachable steady set $\Pi(p_{0})$ is designed to handle precisely these non-ergodic situations: by block decomposition and the data-processing inequality, $V(p_{n+1})\leq V(p_{n})$ always holds; and $V(p_{n})\to 0$ requires, in addition, that each block be primitive (irreducible and aperiodic), yielding a unique invariant $\pi_{j}^{\ast}$. For $p\in\mathcal{P}(\mathcal{X})$, define the block masses
$w_{j}(p)\ :=\ p(\mathcal{X}_{j}),\qquad j=1,\dots,m,$ (6)
and, whenever $w_{j}(p)>0$, the conditional distributions within each block by
$p^{(j)}(x)\ :=\ \begin{cases}\displaystyle\frac{p(x)}{w_{j}(p)},&x\in\mathcal{X}_{j},\\[4.30554pt] 0,&x\notin\mathcal{X}_{j}.\end{cases}$ (7)
Let $T_{j}$ denote the restriction of $T$ to block $\mathcal{X}_{j}$. The set of $T_{j}$-invariant measures is
$\mathcal{I}_{j}\ :=\ \bigl{\{}\ \pi\in\mathcal{P}(\mathcal{X}_{j})\ :\ \pi T_{j}=\pi\ \bigr{\}}.$ (8)
Given an initial distribution $p_{0}$, reachable steady set is defined as
$\Pi(p_{0})\ :=\ \Bigl{\{}\ \pi=\sum_{j=1}^{m}w_{j}(p_{0})\,\pi_{j}\ :\ \pi_{j}\in\mathcal{I}_{j}\ \Bigr{\}}.$ (9)
Intuitively, $\Pi(p_{0})$ collects all steady states that (i) fully relax within each block and (ii) preserve the initial coarse variables $\{w_{j}(p_{0})\}_{j=1}^{m}$ fixed by non-ergodic constraints.
II.2 KLD potential and two basic lemmas
We adopt the Kullback–Leibler divergence (KLD) between $p,q\in\mathcal{P}(\mathcal{X})$,
$D_{\mathrm{KL}}(p\|q)\ :=\ \sum_{x\in\mathcal{X}}p(x)\,\log\!\frac{p(x)}{q(x)},$ (10)
with the convention $0\log(0/q)=0$ and $p\log(p/0)=+\infty$. Given the reachable steady set $\Pi(p_{0})$ in (9), we define KLD potential
$V(p)\ :=\ \inf_{\pi\in\Pi(p_{0})}D_{\mathrm{KL}}(p\|\pi),$ (11)
(cf. the introductory statement (1)). The next two lemmas formalize (i) conservation of block masses and conditional evolution, and (ii) blockwise decomposition of KLD.
Lemma 1 (Preservation of block mass and conditional evolution).
Let $p_{+}=Tp$ with block-invariant $T$ as in (5). Then for all $j=1,\dots,m$,
$w_{j}(p_{+})\ =\ w_{j}(p),$ (12)
and, whenever $w_{j}(p)>0$,
$(Tp)^{(j)}\ =\ T_{j}\,p^{(j)}.$ (13)
Proof.
By block invariance, $T(x,\mathcal{X}_{j})=1$ for $x\in\mathcal{X}_{j}$ and 0 otherwise. Hence $w_{j}(p_{+})=(Tp)(\mathcal{X}_{j})=\sum_{x\in\mathcal{X}}p(x)\,T(x,\mathcal{X}_{j})=\sum_{x\in\mathcal{X}_{j}}p(x)\cdot 1+\sum_{x\notin\mathcal{X}_{j}}p(x)\cdot 0=\sum_{x\in\mathcal{X}_{j}}p(x)=w_{j}(p)$.
which proves (12). For any measurable $A\subseteq\mathcal{X}_{j}$,
$(Tp)^{(j)}(A)=\frac{(Tp)(A)}{w_{j}(p_{+})}=\frac{\sum_{x\in\mathcal{X}}p(x)T(x,A)}{w_{j}(p)}\\ =\sum_{x\in\mathcal{X}_{j}}\tfrac{p(x)}{w_{j}(p)}\,T_{j}(x,A)=(T_{j}p^{(j)})(A).$ (14)
establishing (13). ∎
Lemma 2 (Block decomposition of KLD).
Let $\pi=\sum_{j=1}^{m}w_{j}\,\pi_{j}$ with $w_{j}>0$ and $\pi_{j}\in\mathcal{P}(\mathcal{X}_{j})$ supported on disjoint blocks $\mathcal{X}_{j}$. Then, for any $p\in\mathcal{P}(\mathcal{X})$,
$D_{\mathrm{KL}}(p\|\pi)\ =\ \sum_{j=1}^{m}w_{j}(p)\,D_{\mathrm{KL}}\!\bigl{(}p^{(j)}\|\pi_{j}\bigr{)}\ +\ \sum_{j=1}^{m}w_{j}(p)\,\log\!\frac{w_{j}(p)}{w_{j}}.$ (15)
Proof.
Using the disjointness $\mathcal{X}=\bigsqcup_{j}\mathcal{X}_{j}$ and writing $p(x)=w_{j}(p)\,p^{(j)}(x)$ for $x\in\mathcal{X}_{j}$,
$D_{\mathrm{KL}}(p\|\pi)$ $=\sum_{j=1}^{m}\sum_{x\in\mathcal{X}_{j}}p(x)\,\log\!\frac{p(x)}{w_{j}\,\pi_{j}(x)}$
$=\sum_{j=1}^{m}\sum_{x\in\mathcal{X}_{j}}w_{j}(p)\,p^{(j)}(x)\left[\log\!\frac{p^{(j)}(x)}{\pi_{j}(x)}+\log\!\frac{w_{j}(p)}{w_{j}}\right]$
$=\sum_{j=1}^{m}w_{j}(p)\,D_{\mathrm{KL}}\!\bigl{(}p^{(j)}\|\pi_{j}\bigr{)}+\sum_{j=1}^{m}w_{j}(p)\,\log\!\frac{w_{j}(p)}{w_{j}}.$ (16)
which is (15). ∎
Equations (12)–(13) show that the dynamics preserves coarse variables $\{w_{j}\}$ and evolves conditionals within each block, while (15) cleanly separates within-block divergences from a coarse-mass mismatch term. These identities will be the backbone for the Lyapunov property established in Sec. II.3.
II.3 Lyapunov property of the KLD potential
Motivation and novelty. In many replication scenarios the dynamics is inherently local and respects coarse constraints that are fixed by the substrate: for DNA, heterogeneous base composition (e.g., AT/GC content and sequence-context effects) induces blocks within which substitutions and proofreading occur, while inter-block flow is strongly suppressed. Our novelty is to adopt this block structure explicitly and measure nonequilibrium distance not to a single invariant law, but to the reachable steady set that preserves the initial block masses. Concretely, with the block partition and the reachable set $\Pi(p_{0})$ defined in Sec. II.1, we use the KL potential
$V(p)\ :=\ \inf_{\pi\in\Pi(p_{0})}D_{\mathrm{KL}}(p\|\pi),$
which collapses to the classical KL-to-steady Lyapunov function in the ergodic/single-invariant limit, but remains meaningful under non-ergodicity caused by block invariance.
Statement. We show that $V$ decreases monotonically along any block-invariant Markov evolution $p_{n+1}=Tp_{n}$.
Theorem 1 (KLD potential is Lyapunov under block invariance).
Let $T$ satisfy the block-invariance condition (5) and define $V$ as above. Then, for all $n\geq 0$,
$V(p_{n+1})\ \leq\ V(p_{n}).$ (17)
If, in addition, each block kernel $T_{j}$ is primitive (irreducible and aperiodic) with unique invariant $\pi_{j}^{\ast}\in\mathcal{I}_{j}$, then
$V(p_{n})\ =\ \sum_{j=1}^{m}w_{j}(p_{0})\,D_{\mathrm{KL}}\!\bigl{(}p_{n}^{(j)}\!\|\pi_{j}^{\ast}\bigr{)}\ \xrightarrow[n\to\infty]{}\ 0.$ (18)
Proof.
Fix any $\pi\in\Pi(p_{0})$ and write $\pi=\sum_{j=1}^{m}w_{j}(p_{0})\,\pi_{j}$ with $\pi_{j}\in\mathcal{I}_{j}$. By Lemma 1, $w_{j}(Tp)=w_{j}(p_{0})$ and, by Lemma 2,
$D_{\mathrm{KL}}(Tp\,\|\,\pi)$ $=\sum_{j=1}^{m}w_{j}(p_{0})\,D_{\mathrm{KL}}\!\bigl{(}(Tp)^{(j)}\,\|\,\pi_{j}\bigr{)}$ (19)
$=\sum_{j=1}^{m}w_{j}(p_{0})\,D_{\mathrm{KL}}\!\bigl{(}T_{j}p^{(j)}\,\|\,\pi_{j}\bigr{)}.$
For KLD, the data-processing inequality (DPI) for $f$-divergences gives
$D_{\mathrm{KL}}(T_{j}r\,\|\,\pi_{j})\ \leq\ D_{\mathrm{KL}}(r\,\|\,\pi_{j})\qquad(\pi_{j}T_{j}=\pi_{j}),$ (20)
so inserting (20) into (19) yields
$D_{\mathrm{KL}}(Tp\,\|\,\pi)\ \leq\ \sum_{j=1}^{m}w_{j}(p_{0})\,D_{\mathrm{KL}}\!\bigl{(}p^{(j)}\,\|\,\pi_{j}\bigr{)}\ =\ D_{\mathrm{KL}}(p\,\|\,\pi).$ (21)
Taking the infimum over $\pi\in\Pi(p_{0})$ on both sides proves (17). If each $T_{j}$ is primitive with invariant $\pi_{j}^{\ast}$, then $p_{n}^{(j)}\to\pi_{j}^{\ast}$, and hence (18). ∎
The DPI (hence (20)) is classical; the contribution here is to (i) encode biologically motivated constraints (e.g., heterogeneous AT / GC composition, sequence context locality) as block invariance, (ii) measure distance to a set of steady states that preserves the initial block masses, and (iii) show that the minimal KL to this set remains a Lyapunov function even though the minimizing $\pi$ may change along the trajectory. This covers non-ergodic replication where global mixing is forbidden but within-block relaxation proceeds.
Remark 1 (Equality and strict decay).
Equality in (17) at step $n$ holds iff there exists a minimizer $\pi^{\star}\in\Pi(p_{0})$ with
$D_{\mathrm{KL}}\!\bigl{(}T_{j}p_{n}^{(j)}\,\|\,\pi_{j}^{\star}\bigr{)}=D_{\mathrm{KL}}\!\bigl{(}p_{n}^{(j)}\,\|\,\pi_{j}^{\star}\bigr{)}\quad\text{for all }j,$
i.e., DPI is tight blockwise (this typically means $p_{n}^{(j)}$ is already $T_{j}$-invariant). Otherwise the decrease is strict.
Corollary 1 (Per-step potential drop).
The drop $\Delta V_{n}:=V(p_{n})-V(p_{n+1})\geq 0$ for all $n$, with $\Delta V_{n}=0$ iff the condition in Remark 1 holds.
III Instantiations
III.1 Image replication via Gaussian convolution
We model a single replication step as a smoothing (information loss) Markov map $T_{\sigma}$ induced by discrete Gaussian convolution in an image $I:\mathbb{Z}^{2}\to[0,1]$. Let the normalized Gaussian kernel be
$G_{\sigma}(u,v)\ :=\ \frac{1}{2\pi\sigma^{2}}\,\exp\!\Big{(}-\frac{u^{2}+v^{2}}{2\sigma^{2}}\Big{)},\qquad(u,v)\in\mathbb{Z}^{2},\ \sigma>0,$ (22)
and define the replication map by discrete convolution;
$(I\ast G_{\sigma})(x,y)\ :=\ \sum_{u,v\in\mathbb{Z}}I(x-u,y-v)\,G_{\sigma}(u,v).$ (23)
Let $p_{n}$ denote the (normalized) histogram of pixel intensities before step $n$, viewed as a probability distribution on a finite intensity alphabet $\mathcal{A}$; the one-step channel $T_{\sigma}$ induced by (23) yields
$q_{n}\ =\ p_{n}\,T_{\sigma},\qquad p_{n+1}\ =\ q_{n}.$ (24)
Ergodic vs. non-ergodic (blockwise) blur.
To model non-ergodicity, partition the image grid into disjoint blocks $\{\mathcal{B}_{j}\}_{j=1}^{m}$ and restrict the kernel support so that no mass crosses block boundaries. For $(x,y)\in\mathcal{B}_{j}$,
$(I\ast_{j}G_{\sigma})(x,y)\ :=\ \sum_{\begin{subarray}{c}u,v\in\mathbb{Z}\\ (x-u,y-v)\in\mathcal{B}_{j}\end{subarray}}I(x-u,y-v)\,G_{\sigma}(u,v),$ (25)
which induces a block diagonal channel $T_{\sigma}=\mathrm{diag}\bigl{(}(T_{\sigma})_{1},\dots,(T_{\sigma})_{m}\bigr{)}$ and thus satisfies the block-invariance condition (5). The “ergodic” case corresponds to using (23) globally (no blocking).
Recorded metrics.
At each step we compute (nats)
$H(q_{n})\$$:=\ -\sum_{a\in\mathcal{A}}q_{n}(a)\,\log q_{n}(a),$ (26)
$H_{\times}(n)\$$:=\ H(p_{n},q_{n})\ =\ -\sum_{a\in\mathcal{A}}p_{n}(a)\,\log q_{n}(a),$ (27)
$D_{\mathrm{KL}}(p_{n}\|q_{n})\$$:=\ \sum_{a\in\mathcal{A}}p_{n}(a)\,\log\frac{p_{n}(a)}{q_{n}(a)},$ (28)
$V(p_{n})\$$:=\ \inf_{\pi\in\Pi(p_{0})}D_{\mathrm{KL}}(p_{n}\|\pi),$ (29)
where $\Pi(p_{0})$ is the reachable steady set defined in (9). By Theorem 1, $V(p_{n})$ is nonincreasing along (24) in the non-ergodic (blockwise) case; empirically we also observe monotone $V$ in the ergodic case where the steady set is a singleton.
Figure 1: Ergodic (global) vs. non-ergodic (blockwise) Gaussian blurring on a $256\times 256$ image with $\sigma=1.5$ over $n_{\mathrm{steps}}=50$. Metrics $H$ (26), $H_{\times}$ (27), $D_{\mathrm{KL}}$ (28), and $V$ (29) are reported at each step. In the blockwise case, monotonicity of $V$ follows from Theorem 1.
III.2 DNA replication as a block-diagonal substitution process
Biophysical context.
Single-molecule and theory studies show that DNA polymerases advance stochastically with forward/backward steps, proofreading, and context-dependent kinetics [10, 12, 11, 8, 9]. To capture non-ergodicity induced by compositional structure (e.g. AT- vs. GC-rich segments), we model replication as a block-diagonal Markov process on nucleotides.
Block-diagonal dynamics.
Let $\mathcal{X}=\{\mathrm{A},\mathrm{T},\mathrm{C},\mathrm{G}\}$ with blocks $\mathcal{X}_{1}=\{\mathrm{A},\mathrm{T}\}$ and $\mathcal{X}_{2}=\{\mathrm{C},\mathrm{G}\}$. A one-step substitution kernel (copying) with block invariance (5) is
$T$ $=\operatorname{diag}(T_{1},T_{2}),$ (30)
$T_{1}$ $=\begin{pmatrix}1-\alpha&\alpha\\ \beta&1-\beta\end{pmatrix},\qquad T_{2}=\begin{pmatrix}1-\gamma&\gamma\\ \delta&1-\delta\end{pmatrix}$.
with $\alpha,\beta,\gamma,\delta\in(0,1)$ encoding effective substitution tendencies per step (arising from selectivity, context, and post-replicative processing). The block masses $w_{1}=p_{\mathrm{A}}+p_{\mathrm{T}}$ and $w_{2}=p_{\mathrm{C}}+p_{\mathrm{G}}$ are conserved by the lemma 1.
Proofreading/repair channel.
To incorporate proofreading/repair, we add a blockwise operator $\mathcal{R}=\mathrm{diag}(R_{1},R_{2})$ that is invoked with probability $\rho\in[0,1]$ after extension:
$R_{1}=\begin{pmatrix}1-\alpha^{\prime}&\alpha^{\prime}\\[2.0pt] \beta^{\prime}&1-\beta^{\prime}\end{pmatrix},\qquad R_{2}=\begin{pmatrix}1-\gamma^{\prime}&\gamma^{\prime}\\[2.0pt] \delta^{\prime}&1-\delta^{\prime}\end{pmatrix}$, (31)
with typically $\alpha^{\prime},\beta^{\prime},\gamma^{\prime},\delta^{\prime}<\alpha,\beta,\gamma,\delta$ (improved selectivity). The effective one-step kernel is the convex mixture.
$\widetilde{T}\ :=\ (1-\rho)\,T\ +\ \rho\,\mathcal{R}$, (32)
which remains block diagonal and therefore satisfies (5). The replication update reads
$q_{n}\ =\ p_{n}\,\widetilde{T},\qquad p_{n+1}\ =\ q_{n}.$ (33)
KLD potential under unique block invariants.
If each block chain has a unique invariant $\pi_{j}^{\ast}\in\mathcal{I}_{j}$, then by (11) and (15)
$V(p_{n})\ =\ \sum_{j=1}^{2}w_{j}(p_{0})\,D_{\mathrm{KL}}\!\bigl{(}p_{n}^{(j)}\ \|\ \pi_{j}^{\ast}\bigr{)},$ (34)
and Theorem 1 ensures $V(p_{n+1})\leq V(p_{n})$ and $V(p_{n})\to 0$ as $n\to\infty$.
Energetic bias (phenomenology).
A simple coarse-grained link between proofreading bias and a chemical affinity $\Delta\mu$ reads
$\frac{\alpha^{\prime}}{\beta^{\prime}}\ \approx\ \frac{\alpha}{\beta}\,e^{\Delta\mu/k_{\mathrm{B}}T},\qquad\frac{\gamma^{\prime}}{\delta^{\prime}}\ \approx\ \frac{\gamma}{\delta}\,e^{\Delta\mu/k_{\mathrm{B}}T},$ (35)
suggesting that a stronger driving sharpens $\pi_{j}^{\ast}$ and accelerates the decay of $V$ (cf. kinetic proofreading intuition [5]).
Recorded metrics and bookkeeping.
At each step, we record (in nats): $H(q_{n})$, $H_{\times}(n)=H(p_{n},q_{n})$, $D_{\mathrm{KL}}(p_{n}\|q_{n})$, and $V(p_{n})$.
with monotone $V$ ensured by Theorem 1.
Block-level coarse graining (optional).
Define AT/GC block error rates at step $n$ by
$e_{\mathrm{AT}}(n)\$$:=\ \sum_{x\in\{\mathrm{A},\mathrm{T}\}}\frac{p_{n}(x)}{w_{1}(p_{n})}\sum_{y\in\{\mathrm{C},\mathrm{G}\}}\widetilde{T}_{y|x},$ (36)
$e_{\mathrm{GC}}(n)\$$:=\ \sum_{x\in\{\mathrm{C},\mathrm{G}\}}\frac{p_{n}(x)}{w_{2}(p_{n})}\sum_{y\in\{\mathrm{A},\mathrm{T}\}}\widetilde{T}_{y|x}.$ (37)
The channel induced $2\times 2$ on the coarse alphabet $\{\mathrm{AT},\mathrm{GC}\}$ is
$P(\tilde{y}\mid\tilde{x})\ =\begin{pmatrix}1-e_{\mathrm{AT}}&e_{\mathrm{AT}}\\[2.0pt] e_{\mathrm{GC}}&1-e_{\mathrm{GC}}\end{pmatrix},\qquad\tilde{x},\tilde{y}\in\{\mathrm{AT},\mathrm{GC}\},$ (38)
with input weights $\mathbf{w}(n)=(w_{1}(p_{n}),w_{2}(p_{n}))$. The block-level mutual information (bits) is
$I(\tilde{X};\tilde{Y})\ =\ \sum_{\tilde{x},\tilde{y}}w_{\tilde{x}}(n)\,P(\tilde{y}\mid\tilde{x})\,\log_{2}\!\frac{P(\tilde{y}\mid\tilde{x})}{\sum_{\tilde{x}^{\prime}}w_{\tilde{x}^{\prime}}(n)\,P(\tilde{y}\mid\tilde{x}^{\prime})}$. (39)
Under the symmetric approximation $w_{1}=w_{2}=\tfrac{1}{2}$ and $e_{\mathrm{AT}}=e_{\mathrm{GC}}=\varepsilon$,
[…]
$h_{b}(\varepsilon)$ $=-\varepsilon\log_{2}\varepsilon-(1-\varepsilon)\log_{2}(1-\varepsilon).$
Readouts and figures.
The trajectories in Fig.2 show an increase in $H$ and $H_{\times}$, a decrease in $D_{\mathrm{KL}}(p_{n}\|q_{n})$, and a monotone decay of $V(p_{n})$, in agreement with Theorem 1. Figure 3 shows the potential landscape $V(p)=w_{1}V_{1}(x)+w_{2}V_{2}(y)$ on $(x,y)$, where $x$ is the A fraction in $\{\mathrm{A},\mathrm{T}\}$ and $y$ the C fraction in $\{\mathrm{C},\mathrm{G}\}$.
Figure 2: DNA block-diagonal replication: time series of informational metrics (nats).Shannon entropy, Cross entropy, KLD, and KLD potential. We iterate a four-state Markov chain on $\{\mathrm{A},\mathrm{T},\mathrm{C},\mathrm{G}\}$ with a block-diagonal kernel $\widetilde{T}=\mathrm{diag}(T_{1},T_{2})$ for $N=50$ steps. Base rates: $\alpha=0.020$, $\beta=0.010$, $\gamma=0.015$, $\delta=0.015$. Proofreading channel $\mathcal{R}$ with $\alpha^{\prime}=0.005$, $\beta^{\prime}=0.003$, $\gamma^{\prime}=\delta^{\prime}=0.004$ is mixed at $\rho=0.30$, giving effective rates $\alpha_{\!e}=0.0155$, $\beta_{\!e}=0.0079$, $\gamma_{\!e}=\delta_{\!e}=0.0117$ and block invariants $\pi_{1}^{\ast}=(0.3376,\,0.6624)$, $\pi_{2}^{\ast}=(0.5,\,0.5)$. Initial distribution $p_{0}=(0.6,0.1,0.2,0.1)$ so that $w_{1}(p_{0})=0.7$, $w_{2}(p_{0})=0.3$. At each step $q_{n}=p_{n}\widetilde{T}$ and we record: top-left, Shannon entropy $H(q_{n})$; top-right, cross-entropy $H(p_{n},q_{n})$; bottom-left, step divergence $D_{\mathrm{KL}}(p_{n}\|q_{n})$; bottom-right, KLD potential $V(p_{n})=w_{1}(p_{0})\,D_{\mathrm{KL}}(p_{n}^{(1)}\|\pi_{1}^{\ast})+w_{2}(p_{0})\,D_{\mathrm{KL}}(p_{n}^{(2)}\|\pi_{2}^{\ast})$. As replication proceeds, $H$ and $H_{\times}$ increase, while both $D_{\mathrm{KL}}(p_{n}\|q_{n})$ and $V(p_{n})$ decrease monotonically; the monotonicity of $V$ follows from Theorem 1.
The contour plot shows the KL potential $V(p)=w_{1}D_{\mathrm{KL}}\!\bigl{(}p^{(1)}\!\parallel\!\pi_{1}^{\ast}\bigr{)}+w_{2}D_{\mathrm{KL}}\!\bigl{(}p^{(2)}\!\parallel\!\pi_{2}^{\ast}\bigr{)}$ over $(x,y)\in[0,1]^{2}$, where $x$ is the A fraction in the AT block and $y$ is the C fraction in the GC block. The unique minimum (blue marker) occurs at $(x^{\ast},y^{\ast})=(\pi_{1}^{\ast}(\mathrm{A}),\pi_{2}^{\ast}(\mathrm{C}))\approx(0.338,0.600)$, i.e., at the block-wise invariants determined by the effective rates. Because $V$ is a sum of blockwise KL terms, the level sets are nearly axis-aligned ellipses (here close to circular since $w_{1}=w_{2}=0.5$ and the curvatures are similar). A second-order expansion around $(x^{\ast},y^{\ast})$ gives $V(x,y)\approx\tfrac{w_{1}}{2\,\mu(1-\mu)}(x-\mu)^{2}+\tfrac{w_{2}}{2\,\nu(1-\nu)}(y-\nu)^{2}$ with $\mu=\pi_{1}^{\ast}(\mathrm{A})\approx 0.338$ and $\nu=\pi_{2}^{\ast}(\mathrm{C})=0.600$, yielding coefficients $\simeq 1.12$ and $\simeq 1.04$ (nats), respectively—consistent with the nearly isotropic contours. Increasing proofreading bias decreases $\mu(1-\mu)$ or $\nu(1-\nu)$, steepening the basin and accelerating the monotonic decay of $V(p_{n})$ (Theorem 1). The separable, convex landscape highlights that, under block invariance, the non-ergodic relaxation proceeds independently within blocks and converges to the reachable steady point $(x^{\ast},y^{\ast})$(Fig. 3).
Figure 3: KLD potential landscape $V(p)=w_{1}D_{\mathrm{KL}}(p^{(1)}\parallel\pi_{1}^{\ast})+w_{2}D_{\mathrm{KL}}(p^{(2)}\parallel\pi_{2}^{\ast})$ (nats) over $(x,y)\in[0,1]^{2}$, where $x$ is the A fraction in the AT block and $y$ is the C fraction in the GC block. Proofreading mixture $\rho=0.30$; AT rates $\alpha=0.020$, $\beta=0.010$, $\alpha^{\prime}=0.005$, $\beta^{\prime}=0.003$ give $\alpha_{\mathrm{eff}}=0.0155$, $\beta_{\mathrm{eff}}=0.0079$ and $\pi_{1}^{\ast}=(\beta_{\mathrm{eff}}/(\alpha_{\mathrm{eff}}+\beta_{\mathrm{eff}}),,\alpha_{\mathrm{eff}}/(\alpha_{\mathrm{eff}}+\beta_{\mathrm{eff}}))\approx(0.338,,0.662)$. GC rates (asymmetric) $\gamma=0.014$, $\delta=0.021$, $\gamma^{\prime}=0.004$, $\delta^{\prime}=0.006$ give $\gamma_{\mathrm{eff}}=0.0110$, $\delta_{\mathrm{eff}}=0.0165$ and $\pi_{2}^{\ast}=(\delta_{\mathrm{eff}}/(\gamma_{\mathrm{eff}}+\delta_{\mathrm{eff}}),,\gamma_{\mathrm{eff}}/(\gamma_{\mathrm{eff}}+\delta_{\mathrm{eff}}))\approx(0.60,,0.40)$. Block weights $w_{1}=w_{2}=0.5$. The blue marker indicates $(x^{\ast},y^{\ast})=(\pi_{1}^{\ast}(\mathrm{A}),,\pi_{2}^{\ast}(\mathrm{C}))\approx(0.34,,0.60)$. Grid $101\times 101$, colormap viridis.
An information-theoretic second law (minimal form)
State function and step variables.
Let
$\Delta V_{n}:=V(p_{n})-V(p_{n+1}),$ (41)
and
$D_{n}:=D_{\mathrm{KL}}(p_{n}\|q_{n}),\qquad q_{n}:=p_{n}T.$ (42)
By Theorem 1 and nonnegativity of KLD, one has $\Delta V_{n}\geq 0$ and $D_{n}\geq 0$.
Step production (dimensionless) and its sum.
Define the per-step production
$\mathcal{S}_{n}:=D_{n}+\Delta V_{n},$ (43)
and the cumulative production over $n=0,\dots,N-1$,
$\mathcal{S}^{[0:N)}:=\sum_{n=0}^{N-1}\mathcal{S}_{n}=\sum_{n=0}^{N-1}D_{n}\;+\;V(p_{0})-V(p_{N}).$ (44)
In physical units,
$\sigma_{\rm rep}^{(n)}:=k_{\mathrm{B}}T\,\mathcal{S}_{n},\qquad\Sigma_{\rm rep}^{(N)}:=k_{\mathrm{B}}T\,\mathcal{S}^{[0:N)}.$ (45)
Theorem 2 (Second law for non-ergodic replication: minimal form).
For any block-invariant trajectory $\{p_{n}\}_{n=0}^{N}$,
$\mathcal{S}_{n}\geq 0\quad(n=0,1,\dots,N-1),$ (46)
hence
$\mathcal{S}^{[0:N)}\geq 0.$ (47)
Equivalently, $\sigma_{\rm rep}^{(n)}\geq 0$ and $\Sigma_{\rm rep}^{(N)}\geq 0$.
Idea.
Immediate from $D_{n}\geq 0$ and $\Delta V_{n}\geq 0$. ∎
Remark 2 (Equality and strictness).
$\mathcal{S}_{n}=0$ iff $D_{n}=\Delta V_{n}=0$; otherwise $\mathcal{S}_{n}>0$.
IV Discussion
KLD potential as informational free energy.
Under block invariance, the coarse variables $\{w_{j}\}$ are conserved (Lemma 1) while the within-block conditionals relax toward block-wise invariants. The potential $V$ in (11) therefore measures the nonequilibrium information retained beyond the conserved coarse structure. Its monotone decay (17) provides an information-theoretic second law for non-ergodic replication: fine-scale distinguishability is irreversibly lost, whereas coarse composition is preserved.
Per-step bookkeeping and physical units.
The minimal second law above decomposes the step production into two nonnegative parts, $D_{n}$ and $\Delta V_{n}$, cf. (43). In physical units, the potential drop defines an informational free-energy change
$\Delta F_{\mathrm{info}}^{(n)}:=k_{\mathrm{B}}T\,\Delta V_{n}=k_{\mathrm{B}}T\,[\,V(p_{n})-V(p_{n+1})\,]\ \geq 0,$ (48)
while $k_{\mathrm{B}}T\,D_{n}$ quantifies the stepwise dissipated “informational heat.” Together they account for the net degradation incurred by one replication step.
Ergodic vs. non-ergodic limits.
If the reachable steady set $\Pi(p_{0})$ collapses to a single invariant distribution (ergodic limit), $V$ reduces to the standard KL-to-steady Lyapunov function. Under non-ergodic, block-invariant constraints, $V$ is the minimal KL distance to the steady set $\Pi(p_{0})$, so its decay quantifies relaxation under conserved coarse masses. This distinction is essential for replication processes where structural constraints inhibit global mixing.
Tightness and equality cases.
Equality in (17) requires blockwise tightness of the data-processing inequality (Remark 1); operationally, this corresponds to being at (or effectively on) a block-wise steady state. Away from such fixed points, the decrease is strict and $V$ serves as a sensitive progress variable for replication-driven relaxation.
V Conclusion and Outlook
We formulate replication as a discrete Markov map under block invariance and prove that minimal KLD to the reachable steady set,
$V(p)=\inf_{\pi\in\Pi(p_{0})}D_{\mathrm{KL}}(p\|\pi),$
is a Lyapunov function (Theorem 1). This yields an information-theoretic second law for non-ergodic replication, complementary to Landauer-type results for erasure. Instantiations in Gaussian image copying and DNA block-diagonal substitution with proofreading exhibit the predicted behavior: monotone $V$, increasing $H$ and cross-entropy, and decreasing $D_{\mathrm{KL}}(p_{n}\|q_{n})$. Future directions include (i) force-resolved single-molecule tests of potential drops and no-free-copying bounds, (ii) extensions to multiscale block hierarchies and heterogeneous kinetics, and (iii) quantum/continuous-state analogues where contractive metrics (e.g. quantum $f$-divergences) may provide replication second laws under conserved coarse observables.
Appendix A Simulation protocol for Fig. 1 (Gaussian-copy model)
Software and environment. All results in Fig. 1 were generated in Python (NumPy/SciPy/Matplotlib). Gaussian smoothing uses scipy.ndimage.gaussian_filter with boundary condition mode="reflect".
Domain and initialization. We use a $256\times 256$ pixel grid. The image domain is partitioned into a regular $B_{x}\times B_{y}$ tiling of rectangular blocks (Fig. 1 uses $B_{x}=B_{y}=4$; we also verified $2\times 2$ and $128\times 128$). Inside each block, a binary pattern is drawn i.i.d. from a Bernoulli law with a checkerboard success probability: even (i+j)(i{+}j) blocks use $0.8$, odd (i+j)(i{+}j) blocks use $0.2$. The random seed is fixed to $7$. The resulting array is normalized to a probability mass function (pmf) $p_{0}=I_{0}/\!\sum_{x,y}I_{0}(x,y)$.
Replication step ($n\!\to\!n{+}1$). Let $p_{n}$ denote the current pmf on pixels. We set $\sigma=1.5$ pixels and apply a Gaussian smoothing, followed by renormalization so that $\sum_{x,y}q_{n}(x,y)=1$. Two cases are considered:
  • Ergodic (global): Gaussian smoothing is applied to the whole image to produce $q_{n}$.
  • Non-ergodic (blockwise): The image is split into blocks; the same Gaussian is applied independently within each block (no cross-block smoothing). The blocks are then stitched back together to form $q_{n}$. This preserves the total mass inside each block at every step.
In both cases we update $p_{n+1}=q_{n}$. We run $n_{\text{steps}}=50$ and store snapshots at $n\in\{0,10,20,30,40,50\}$.
Information measures (base =2=2, in bits). At each step we compute
$H(q_{n})$ $:=-\sum_{x,y}q_{n}\log_{2}q_{n},$ (49)
$H(p_{n},q_{n})$ $:=-\sum_{x,y}p_{n}\log_{2}q_{n},$ (50)
$D_{\mathrm{KL}}(p_{n}\|q_{n})$ $:=\sum_{x,y}p_{n}\log_{2}\!\frac{p_{n}}{q_{n}},$ (51)
where terms with $p_{n}=0$ or $q_{n}=0$ are omitted in the sum to avoid $\log 0$. (To convert bits to nats multiply by $\ln 2$.)
KLD potential $V(p_{n})$. We use the Lyapunov-type potential defined consistently with the main text:
  • Ergodic case: $V(p_{n})=D_{\mathrm{KL}}(p_{n}\|u)$, where $u$ is the uniform pmf over all $256\times 256$ pixels.
  • Non-ergodic (blockwise) case: Let $b$ index blocks, $w_{b}=\sum_{(x,y)\in b}p_{n}(x,y)$ the block mass, $p_{b}$ the within-block normalized pmf, and $u_{b}$ the blockwise uniform pmf. Then
    $V(p_{n})=\sum_{b}w_{b}\,D_{\mathrm{KL}}\!\bigl{(}p_{b}\|u_{b}\bigr{)}.$ (52)
This definition matches the theoretical reachable-steady-set construction and yields the monotone decay predicted by Theorem 1.
Visualization and units. Time-series plots show $H$, $H(p_{n},q_{n})$, $D_{\mathrm{KL}}(p_{n}\|q_{n})$, and $V(p_{n})$ versus the replication step $n$. Snapshots at the specified steps use a common grayscale colorbar. All values are reported in bits.
Sanity condition on $\sigma$. If the Gaussian width exceeds roughly one quarter of the smallest block edge, boundary effects may dominate. In Fig. 1 (block size $64{\times}64$, $\sigma=1.5$) this issue does not arise. The same qualitative trends were observed for $2{\times}2$ and $128{\times}128$ partitions (not shown).
Appendix B: Simulation details for Fig. 2 (time series)
B.1 Model and parameters.
We consider a block-diagonal, discrete-time Markov kernel on $\mathcal{X}=\{\mathrm{A},\mathrm{T},\mathrm{C},\mathrm{G}\}$ with blocks $\mathcal{X}_{1}=\{\mathrm{A},\mathrm{T}\}$ and $\mathcal{X}_{2}=\{\mathrm{C},\mathrm{G}\}$. One replication step uses an extension channel $T$ mixed with a block-wise proofreading/repair channel $\mathcal{R}$ with probability $\rho$:
$\widetilde{T}\;=\;(1-\rho)\,T\;+\;\rho\,\mathcal{R}\;=\;\mathrm{diag}\!\bigl{(}T_{1}^{\mathrm{eff}},T_{2}^{\mathrm{eff}}\bigr{)}.$
Each $2\times 2$ block takes the form
$T_{1}=\begin{pmatrix}1-\alpha&\alpha\\[2.0pt] \beta&1-\beta\end{pmatrix},\quad T_{2}=\begin{pmatrix}1-\gamma&\gamma\\[2.0pt] \delta&1-\delta\end{pmatrix}$,
$R_{1}=\begin{pmatrix}1-\alpha^{\prime}&\alpha^{\prime}\\[2.0pt] \beta^{\prime}&1-\beta^{\prime}\end{pmatrix},\quad R_{2}=\begin{pmatrix}1-\gamma^{\prime}&\gamma^{\prime}\\[2.0pt] \delta^{\prime}&1-\delta^{\prime}\end{pmatrix}$.
Effective (mixed) rates are
$\alpha_{\!\mathrm{eff}}=(1-\rho)\alpha+\rho\alpha^{\prime},\ \beta_{\!\mathrm{eff}}=(1-\rho)\beta+\rho\beta^{\prime},$
$\gamma_{\!\mathrm{eff}}=(1-\rho)\gamma+\rho\gamma^{\prime},\ \delta_{\!\mathrm{eff}}=(1-\rho)\delta+\rho\delta^{\prime}.$
Thus
$T_{1}^{\mathrm{eff}}=\begin{pmatrix}1-\alpha_{\!\mathrm{eff}}&\alpha_{\!\mathrm{eff}}\\[2.0pt] \beta_{\!\mathrm{eff}}&1-\beta_{\!\mathrm{eff}}\end{pmatrix},\quad T_{2}^{\mathrm{eff}}=\begin{pmatrix}1-\gamma_{\!\mathrm{eff}}&\gamma_{\!\mathrm{eff}}\\[2.0pt] \delta_{\!\mathrm{eff}}&1-\delta_{\!\mathrm{eff}}\end{pmatrix}$.
Block-wise invariants are
$\pi_{1}^{\ast}=\!\Bigl{(}\tfrac{\beta_{\!\mathrm{eff}}}{\alpha_{\!\mathrm{eff}}+\beta_{\!\mathrm{eff}}},\tfrac{\alpha_{\!\mathrm{eff}}}{\alpha_{\!\mathrm{eff}}+\beta_{\!\mathrm{eff}}}\Bigr{)}\approx(0.3376,\,0.6624),$
$\pi_{2}^{\ast}=\!\Bigl{(}\tfrac{\delta_{\!\mathrm{eff}}}{\gamma_{\!\mathrm{eff}}+\delta_{\!\mathrm{eff}}},\tfrac{\gamma_{\!\mathrm{eff}}}{\gamma_{\!\mathrm{eff}}+\delta_{\!\mathrm{eff}}}\Bigr{)}=(0.5,\,0.5).$
B.2 Initial condition and conserved masses.
We initialize $p_{0}=(p_{A},p_{T},p_{C},p_{G})=(0.6,0.1,0.2,0.1)$ and normalize. Block masses
$w_{1}(p):=p_{A}+p_{T},\qquad w_{2}(p):=p_{C}+p_{G}$
are conserved by block invariance; we store $w_{1}(p_{0})$ and $w_{2}(p_{0})$ for the potential below.
B.3 Recorded quantities (natural logs; nats).
At step $n$ we set $q_{n}:=p_{n}\,\widetilde{T}$ (for $n<N$) and record
$H(q_{n})=-\!\sum_{x}q_{n}(x)\log q_{n}(x),\quad H_{\times}(n)=-\!\sum_{x}p_{n}(x)\log q_{n}(x),$
$D_{\mathrm{KL}}(p_{n}\|q_{n})=\sum_{x}p_{n}(x)\log\!\frac{p_{n}(x)}{q_{n}(x)}.$
The KLD potential (Def. (11)) reduces to the block sum
$V(p_{n})=w_{1}(p_{0})\,D_{\mathrm{KL}}\!\bigl{(}p_{n}^{(1)}\|\pi_{1}^{\ast}\bigr{)}+w_{2}(p_{0})\,D_{\mathrm{KL}}\!\bigl{(}p_{n}^{(2)}\|\pi_{2}^{\ast}\bigr{)},$
with $p_{n}^{(1)}=(p_{A},p_{T})/w_{1}(p_{n})$ and $p_{n}^{(2)}=(p_{C},p_{G})/w_{2}(p_{n})$.
B.4 Iteration.
We iterate
$p_{n+1}=q_{n}=p_{n}\,\widetilde{T},\qquad n=0,1,\dots,N-1,$
with $N=50$. At each $n$ we first record $V(p_{n})$; then, if $n<N$, we compute and record $H(q_{n})$, $H_{\times}(n)$, and $D_{\mathrm{KL}}(p_{n}\|q_{n})$, and finally set $p_{n+1}=q_{n}$.
B.5 Minimal Python excerpt (reproducibility).
import numpy as np
def H_shannon(p):
    m = p > 0
    return -np.sum(p[m]*np.log(p[m]))
def H_cross(p, q):
    m = (p > 0) & (q > 0)
    return -np.sum(p[m]*np.log(q[m]))
def Dkl(p, q):
    m = (p > 0) & (q > 0)
    return np.sum(p[m]*(np.log(p[m]) - np.log(q[m])))
def V_block2(p, w1_0, w2_0, pi1, pi2):
    pA,pT,pC,pG = p
    w1, w2 = pA+pT, pC+pG
    p1 = np.array([pA, pT])/w1
    p2 = np.array([pC, pG])/w2
    return w1_0*Dkl(p1, pi1) + w2_0*Dkl(p2, pi2)
# parameters (nats)
alpha, beta   = 0.020, 0.010
gamma, delta  = 0.015, 0.015
alpha_p, beta_p = 0.005, 0.003
gamma_p, delta_p = 0.004, 0.004
rho = 0.30
ae = (1-rho)*alpha + rho*alpha_p
be = (1-rho)*beta  + rho*beta_p
ge = (1-rho)*gamma + rho*gamma_p
de = (1-rho)*delta + rho*delta_p
T1 = np.array([[1-ae, ae],[be, 1-be]], float)
T2 = np.array([[1-ge, ge],[de, 1-de]], float)
T  = np.zeros((4,4), float); T[:2,:2]=T1; T[2:,2:]=T2
pi1 = np.array([be/(ae+be), ae/(ae+be)], float)
pi2 = np.array([de/(ge+de), ge/(ge+de)], float)
p = np.array([0.6,0.1,0.2,0.1], float); p = p/np.sum(p)
w1_0, w2_0 = p[0]+p[1], p[2]+p[3]
N = 50
Hq, Hx, D, V = [], [], [], []
for n in range(N+1):
    V.append(V_block2(p, w1_0, w2_0, pi1, pi2))   # V(p_n)
    if n == N: break
    q = p @ T # q_n = p_n T
    Hq.append(H_shannon(q))                       # H(q_n)
    Hx.append(H_cross(p, q))                      # H(p_n, q_n)
    D.append(Dkl(p, q))                           # D_KL(p_n || q_n)
    p = q
Appendix C: Simulation details for Fig. 3 (potential landscape)
C.1 Model and objective.
We consider two invariant blocks $\mathcal{X}_{1}=\{\mathrm{A},\mathrm{T}\}$ and $\mathcal{X}_{2}=\{\mathrm{C},\mathrm{G}\}$. On the grid $(x,y)\in[0,1]^{2}$, we parameterize a block-wise distribution
$p(x,y)=\bigl{(}w_{1}x,\;w_{1}(1-x),\;w_{2}y,\;w_{2}(1-y)\bigr{)},$
where $x$ is the A-fraction in the AT block and $y$ is the C-fraction in the GC block. The KL potential (Def. (11)) reduces to the separable sum
$V\bigl{(}p(x,y)\bigr{)}=w_{1}\,D_{\mathrm{KL}}\!\bigl{(}[x,1-x]\|\pi_{1}^{\ast}\bigr{)}\;+\;w_{2}\,D_{\mathrm{KL}}\!\bigl{(}[y,1-y]\|\pi_{2}^{\ast}\bigr{)},$
in nats. We plot $V$ over $0,1]^{2}$ and mark the minimum $(x^{\ast},y^{\ast})=\bigl{(}\pi_{1}^{\ast}(\mathrm{A}),\,\pi_{2}^{\ast}(\mathrm{C})\bigr{)}$.
C.2 Effective rates and steady states.
One step is a convex mixture of an extension channel $T$ and a proofreading channel $\mathcal{R}$ with probability $\rho=0.30$. Each $2\times 2$ block is
$T_{1}=\begin{pmatrix}1-\alpha&\alpha\\ \beta&1-\beta\end{pmatrix},\quad R_{1}=\begin{pmatrix}1-\alpha^{\prime}&\alpha^{\prime}\\ \beta^{\prime}&1-\beta^{\prime}\end{pmatrix}$,
$T_{2}=\begin{pmatrix}1-\gamma&\gamma\\ \delta&1-\delta\end{pmatrix},\quad R_{2}=\begin{pmatrix}1-\gamma^{\prime}&\gamma^{\prime}\\ \delta^{\prime}&1-\delta^{\prime}\end{pmatrix}$,
so the effective rates are
$\alpha_{\!e}$ $=(1-\rho)\alpha+\rho\alpha^{\prime},$ $\beta_{\!e}$ $=(1-\rho)\beta+\rho\beta^{\prime},$ (53)
$\gamma_{\!e}$ $=(1-\rho)\gamma+\rho\gamma^{\prime},$ $\delta_{\!e}$ $=(1-\rho)\delta+\rho\delta^{\prime}.$
AT block (Fig. 2): $\alpha=0.020,\ \beta=0.010,\ \alpha^{\prime}=0.005,\ \beta^{\prime}=0.003$ $\Rightarrow\ \alpha_{\!e}=0.0155,\ \beta_{\!e}=0.0079$. GC block: $\gamma=0.014,\ \delta=0.021,\ \gamma^{\prime}=0.004,\ \delta^{\prime}=0.006$ $\Rightarrow\ \gamma_{\!e}=0.0110,\ \delta_{\!e}=0.0165$. The block-wise invariants (for [[1a,a],[b,1b]][[1-a,a],[b,1-b]]) are
$\pi_{1}^{\ast}=\Bigl{(}\tfrac{\beta_{\!e}}{\alpha_{\!e}+\beta_{\!e}},\;\tfrac{\alpha_{\!e}}{\alpha_{\!e}+\beta_{\!e}}\Bigr{)}\approx(0.338,\,0.662),$
$\pi_{2}^{\ast}=\Bigl{(}\tfrac{\delta_{\!e}}{\gamma_{\!e}+\delta_{\!e}},\;\tfrac{\gamma_{\!e}}{\gamma_{\!e}+\delta_{\!e}}\Bigr{)}=(0.60,\,0.40).$
We take $(w_{1},w_{2})=(0.5,0.5)$ unless otherwise noted. The blue marker in Fig. 3 is placed at $(x^{\ast},y^{\ast})=(\pi_{1}^{\ast}(\mathrm{A}),\pi_{2}^{\ast}(\mathrm{C}))\approx(0.338,0.600)$.
C.3 Grid and evaluation.
We use a $101\times 101$ grid on $0,1]^{2}$. For Bernoulli pairs we evaluate
$D_{\mathrm{KL}}\!\bigl{(}[p,1-p]\|[q,1-q]\bigr{)}=p\log\tfrac{p}{q}+(1-p)\log\tfrac{1-p}{1-q},$
with safe clipping to avoid $\log 0$. The color scale shows $V$ in nats (“viridis” colormap). Axes: $x$ = A fraction (AT), $y$ = C fraction (GC).
C.4 Minimal Python excerpt (reproducibility).
import numpy as np
# mixture rate
rho = 0.30
# AT block (same as Fig.2)
alpha, beta   = 0.020, 0.010
alpha_p, beta_p = 0.005, 0.003
ae = (1 - rho)*alpha + rho*alpha_p      # 0.0155
be = (1 - rho)*beta  + rho*beta_p       # 0.0079
# GC block (asymmetric, to place y*  0.60)
gamma, delta   = 0.014, 0.021
gamma_p, delta_p = 0.004, 0.006
ge = (1 - rho)*gamma + rho*gamma_p      # 0.0110
de = (1 - rho)*delta + rho*delta_p      # 0.0165
# block weights
w1, w2 = 0.5, 0.5
# invariants for [[1-a, a],[b, 1-b]]
pi1_A = be/(ae+be);  pi1_T = ae/(ae+be)     #  (0.338, 0.662)
pi2_C = de/(ge+de);  pi2_G = ge/(ge+de)     #   (0.600, 0.400)
def kl_bernoulli(p, q, eps=1e-12):
    p = np.clip(p, eps, 1.0 - eps)
    q = np.clip(q, eps, 1.0 - eps)
    return p*np.log(p/q) + (1-p)*np.log((1-p)/(1-q))
# grid and potential values (nats)
N = 101
xs = np.linspace(0.0, 1.0, N)   # A-fraction in AT
ys = np.linspace(0.0, 1.0, N)   # C-fraction in GC
X, Y = np.meshgrid(xs, ys, indexing=’xy’)
V = w1*kl_bernoulli(X, pi1_A) + w2*kl_bernoulli(Y, pi2_C)
# minimum location for the blue marker:
x_star, y_star = pi1_A, pi2_C
This excerpt computes the potential field $V(p)$ on the grid and the minimum $(x^{\ast},y^{\ast})$. The plotted Fig. 3 uses a filled contour of $V$, the “viridis” colormap, and annotates $(x^{\ast},y^{\ast})$ with a blue marker.